Is the following sequence arithmetic or geometric?
1/2, 1/10, 1/50, 1/250, 1/1250...
Find the common difference or ratio depending on your decision to the first part of the question.

Answer by RAY100(1637)

Geometric with a common ratio of (1/5)
.
1/2 *1/5 =1/10,,*1/5 = 1/50,,*1/5 =1/250,,*1/5=1/1250,,